# Approximation in Hölder spaces

### v3

@@ 1.5. Applications to compactness of the bi-commutator
where $T_{i}$ acts on the variable $x_{i}.$ The product BMO upper bound for the $L^{p}\to L^{p}$ boundedness of the bi-commutator was first obtained with the Hilbert transforms in Ferguson, Sadosky [FerSad], and in full generality of completely arbitrary CZOs later by Dalenc, Ou [DaO16]. The corresponding lower bound in terms of product $\mathrm{BMO}$ for special singular integrals (Hilbert/Riesz transforms) has a somewhat complicated history. While the product $\mathrm{BMO}$ is still believed to be the characterizing condition, the existing proofs have been reported to have a gap, see e.g. [Ferguson2002, HTV2021ce, AHLMO2021], and to the best of our knowledge a fix has not yet been published. The existence of this gap adds considerable interest to all natural questions involving the bi-commutator.
@@
Now, in place of considering $L^{p}$-to-$L^{p}$ bounds it is natural to consider mixed $L^{p_{1}}(L^{p_{2}})$-to-$L^{q_{1}}(L^{q_{2}})$ bounds, where we allow $q_{i}\not=p_{i},$ this was done recently in [AHLMO2021, Theorem 1.1.]. The exponents that directly connect with our work are those for which $\beta_{i}:=n_{i}(1/p_{i}-1/q_{i})\geq 0$ and $\beta_{i}>0$ for some $i=1,2,$ which we assume here and below. Then, the bi-parameter space to characterize mixed $L^{p_{1}}(L^{p_{2}})$-to-$L^{q_{1}}(L^{q_{2}})$ boundedness turned out to be
$\|b$ $\|_{\operatorname{BMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n})}=\sup_{R=I_{1}\times I_{2}}\mathcal{O}^{\beta_{1},\beta_{2}}(b,I_{1}\times I_{2})$
@@
$=\sup_{R=I_{1}\times I_{2}}\ell(I_{1})^{-\beta_{1}}\fint_{I_{1}}\ell(I_{2})^{-\beta_{2}}\fint_{I_{2}}|b(x_{1},x_{2})-\langle b(x_{1},\cdot)\rangle_{I_{2}}-\langle b(\cdot,x_{2})\rangle_{I_{1}}+\langle b\rangle_{I_{1}\times I_{2}}|\,\mathrm{d}x_{2}\,\mathrm{d}x_{1}.$
@@
The space $\operatorname{BMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n})$ is in fact a $\mathrm{BMO}$-valued or $\dot{C}^{0,\alpha}$-valued Hölder space , see [AHLMO2021, Section 3]. Then, very recently [MarOik24, Theorem 1.1.] the corresponding $L^{p_{1}}(L^{p_{2}})$-to-$L^{q_{1}}(L^{q_{2}})$ compactness was characterized through the VMO analogue $\operatorname{VMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n}).$ Moreover, as a corollary of these truly mixed norm compactness characterizations the natural product BMO condition in the non-mixed $L^{p}\to L^{p}$ case was attained as a sufficient condition. One of the crucial theorems in [MarOik24] is the following theorem, which is, in fact, a corollary of the approximations developed in this article.
@@
Let $\beta_{i}\in[0,1)$ and $\beta_{i}>0$ for some $i=1,2,$ then
@@
Theorem 1.11 is a bi-parameter analogue of Theorem 1.10 (with $Y=\mathbb{C}$), where in at least one parameter we have a vanishing Hölder type behaviour. The proof of Theorem 1.11 is split between the current article and [MarOik24]. Indeed, we record here a quantitative version of Theorem 1.6 in the body-text after the proof of Theorem 1.6 as Corollary 2.5; then, assuming Corollary 2.5, Theorem 1.11 is proved in [MarOik24, Section 3.].
@@ Acknowledgements
C. M. was supported by the grant no. 334466 of the Research Council of Norway, “Fourier Methods and Multiplicative Analysis”. T.O. was supported by the Finnish Academy of Science and Letters, and by the MICINN (Spain) grant no. PID2020-114167GB-I00.
@@ 2. Bounded support approximation
In this section we prove Theorem 1.2 and then show how to obtain from this Theorem 1.6 and then Corollary 2.5. After these, at the end of this section, we prove Theorem 1.9. Recall that we simplify notation by denoting $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|.$ We begin with a simple remark.
@@
First, note that if $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and has bounded support, then, by property (3) of $\omega,$ $f$ belongs to $\dot{\operatorname{VC}}^{0,\omega}(X,Y).$ Since $\dot{\operatorname{VC}}^{0,\omega}(X,Y)$ is closed under limits with respect to the seminorm $\dot{C}^{0,\omega},$ as per Remark 2.1, the inclusion ``"``"`$\subset$"` ` ⊂ " holds true.
To prove the reverse (much more complicated) inclusion, let $f\in\dot{\operatorname{VC}}^{0,\omega}$ and $\varepsilon>0.$ By Lemma 2.2 we find two parameters $0<r\ll R$ such that
$\sup_{\begin{subarray}{c}x\not=y\in X\\ \|x-y\|\leq r\end{subarray}}\frac{\left|f(x)-f(y)\right|}{\omega(\|x-y\|)}\leq\varepsilon,$ (12)
and
@@
We consider an auxiliary parameter $M\gg R$ and the following function
$\tau=\tau_{M}:X\to B(0,M),\qquad\tau(x)=\begin{cases}x,\qquad&\|x\|<M,\\ \left(\frac{2M-\|x\|}{M}\right)^{2}x,\qquad M\leq&\|x\|<2M,\\ 0,\qquad&\|x\|\geq 2M.\end{cases}$ (14)
We next show that if $r$ is small enough and $R,M$ are large enough then
$f\circ\tau\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}},\qquad\|f-f\circ\tau\|_{\dot{C}^{0,\omega}}\lesssim\varepsilon.$ (15)
Notice that $f\circ\tau-f(0)$ is zero outside $B(0,2M),$ and hence $f\circ\tau$ has bounded support. (Recall that $\dot{C}^{0,\omega}$ is defined modulo constant equivalence classes.) To actually verify the claims of (15) we begin checking some properties of the function $\tau.$
First, we verify
$\|\tau(x)-\tau(z)\|\leq 5\|x-z\|\quad\text{for all}\quad x,z\in X,$ (16)
and in particular by doing so, Lemma 2.3 implies the left claim on the line (15). Indeed, for points $x,z\in B(0,2M)\setminus B(0,M),$ the definition of $\tau$ and the triangle inequality give
$\begin{split}\|\tau(x)-\tau(z)\|&\leq\Big{|}\big{(}2-\frac{\|x\|}{M}\big{)}^{2}-\big{(}2-\frac{\|z\|}{M}\big{)}^{2}\Big{|}\|z\|+\big{(}2-\frac{\|x\|}{M}\big{)}^{2}\|x-z\|\\ &\leq\big{(}2-\frac{\|x\|}{M}+2-\frac{\|z\|}{M}\big{)}\frac{\|x-z\|}{M}\|z\|+\big{(}2-\frac{\|x\|}{M}\big{)}^{2}\|x-z\|\\ &\leq\big{(}4-\frac{2M}{M}\big{)}\frac{\|x-z\|}{M}(2M)+\big{(}2-\frac{M}{M}\big{)}^{2}\|x-z\|=5\|x-z\|,\end{split}$ (17)
where in the second bound we used the basic identity $a^{2}-b^{2}=(a-b)(a+b).$ As $\tau$ is obviously $1$-Lipschitz in the sets $B(0,M)$ and $B(0,2M)^{c},$ and $\tau$ is continuous in $X,$ we deduce (16) for all $x,z\in X.$
Then we show the following contraction property:
$\begin{split}\Big{\{}x\in B(0,M)^{c},\quad\tau(x),\tau(z)\in B(0,R)\Big{\}}\Longrightarrow\|\tau(x)-\tau(z)\|\lesssim R\,\theta(M)\|x-z\|,\end{split}$ (18)
where $\theta:(0,\infty)\to(0,\infty)$ is a function satisfying $\lim_{M\to\infty}\theta(M)=0$, whose definition will be specified while we next check the truth of the implication [^1] We remark that (18) would not be true, if we replaced the power $2$ in the definition (14) of $\tau$ by the power $1$. If $\|x-z\|>M/2,$ then we have
$\|\tau(x)-\tau(z)\|\leq 2R<2R\frac{\|x-z\|}{M/2}=4R\,\theta_{1}(M)\|x-z\|,\qquad\theta_{1}(M):=\frac{1}{M}.$
So next assume that $z\in B(x,M/2).$ Since $R\ll M$ and $\tau(z)\in B(0,R)$, by the definition of $\tau$ we must have necessarily $\|z\|>M.$ Thus, it remains to show that
$\begin{split}\Big{\{}x,z\in B(0,M)^{c},\quad\tau(x),\tau(z)\in B(0,R)\Big{\}}\Longrightarrow\|\tau(x)-\tau(z)\|\lesssim R\,\theta(M)\|x-z\|,\end{split}$ (19)
which is symmetric with respect to both variables $x,z.$ Notice that if $x,z\in B(0,2M)^{c},$ then the left-hand side of the claimed estimate is zero. Assume that $x,z\in B(0,2M)\setminus B(0,M)$ and we make the following observation. By $\tau(x)\in B(0,R)$ and $x\in B(0,M)^{c},$ there holds that
$\Big{\|}\left(\frac{2M-\|x\|}{M}\right)^{2}x\Big{\|}\leq R\Longrightarrow\Big{|}\frac{2M-\|x\|}{M}\Big{|}\leq\sqrt{\frac{R}{\|x\|}}\leq\sqrt{\frac{R}{M}}=\sqrt{R}\,\theta_{2}(M),\qquad\theta_{2}(M):=\frac{1}{\sqrt{M}},$
and the same bound is valid for the variable $z.$ Thus, by (17) we obtain
$\|\tau(x)-\tau(z)\|$ $\leq\Big{(}2-\frac{\|x\|}{M}+2-\frac{\|z\|}{M}\Big{)}\frac{\|x-z\|}{M}\|z\|+\Big{(}2-\frac{\|x\|}{M}\Big{)}^{2}\|x-z\|$
$\leq\Big{(}4\sqrt{\frac{R}{M}}+\frac{R}{M}\Big{)}\|x-z\|\leq 5R\,\theta_{2}(M)\|x-z\|.$
As the last case, suppose that $z\in B(0,2M)^{c}.$ This case follows by the continuity of $\tau$ and the previous case. Indeed, taking a point $y\in[x,z]$ with $\|y\|=2M,$ by the previous case applied for $x$ and $y,$ and bearing in mind that $\tau(z)=\tau(y)=0,$ we deduce
$\|\tau(x)-\tau(z)\|=\|\tau(x)-\tau(y)\|\leq 5R\,\theta_{2}(M)\|x-y\|\leq 5R\,\theta_{2}(M)\|x-z\|.$
All in all, we have now shown that (18) is valid with the function $\theta:(0,\infty)\to(0,\infty)$ given by
$\theta(M):=\frac{1}{\sqrt{M}}\geq\frac{1}{2}(\theta_{1}(M)+\theta_{2}(M)).$ (20)
Now, let us denote $g:=f\circ\tau\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}$ and let $r^{\prime}>0$ be such that
$\|g\|_{\dot{C}^{0,\omega}_{r^{\prime}}}:=\sup_{\begin{subarray}{c}x\not=y\in X\\ \|x-y\|<r^{\prime}\end{subarray}}\frac{|g(x)-g(y)|}{\omega(\|x-y\|)}\leq\varepsilon.$ (21)
By (12), the fact that $\operatorname{Lip}(\tau)\leq 5$ (see (16)), and the proof of Lemma 2.3, this new $r^{\prime}>0$ can be taken to be an absolute multiple of the $r$ in (12). Therefore, we may and do assume that both (12) and (21) hold for the same $r=r(\varepsilon),$ which is of course independent of $R$ and $M.$
Next we show that
$\sup_{x\in B(0,M)^{c}}\sup_{z\in X}\frac{\left|g(x)-g(z)\right|}{\omega(\|x-z\|)}\lesssim\varepsilon,$ (22)
provided $M$ is large enough. By (21), it is enough to estimate (22) only at those couples $x,z$ with $\|x-z\|\geq r.$ Fix a point $x\in B(0,M)^{c}$ and we distinguish into cases. If $\tau(x)\in B(0,R)^{c},$ then by (13) and $\omega$ being non-decreasing and doubling we obtain
$\begin{split}\left|g(x)-g(z)\right|&=\left|f(\tau(x))-f(\tau(z))\right|\\ &\leq\Big{(}\sup_{u\in B(0,R)^{c}}\sup_{v\in X}\frac{\left|f(u)-f(v)\right|}{\omega(\|u-v\|)}\Big{)}\omega(\|\tau(x)-\tau(z)\|)\lesssim_{C_{\operatorname{db}}}\varepsilon\omega(\|x-z\|).\end{split}$ (23)
The case $\tau(z)\in B(0,R)^{c}$ is symmetrical.
Now consider $\tau(x),\tau(z)\in B(0,R)$ and we split into subcases. But first observe that for those $x,z$ with $\|x-z\|\geq M^{1/4},$ using that $f\in\dot{C}^{0,\omega}$ we can write
$\left|g(x)-g(z)\right|=\left|f(\tau(x))-f(\tau(z))\right|\leq\|f\|_{\dot{C}^{0,\omega}}\omega(|\tau(x)-\tau(z)|)\leq\|f\|_{\dot{C}^{0,\omega}}\omega(2R).$
Provided that $M=M(\varepsilon,R)$ is large enough, the condition $\omega(\infty)=\infty$ says that the last term can be made smaller than $\varepsilon\omega(M^{1/4})\leq\varepsilon\omega(\|x-z\|),$ and thus (22) holds for these $x$ and $z.$ Therefore, during the rest of the verification of (22) we can assume that $\|x-z\|\leq M^{1/4},$ whenever needed.
Notice also that in the case $x,z\in X\setminus B(0,2M)$ there is nothing to prove, as then $g(x)=g(z)=0.$
Thus, we can assume that both $x,z\in B(0,10M).$ In this case we bound
$\begin{split}\left|g(x)-g(z)\right|&=\left|f(\tau(x))-f(\tau(z))\right|\\ &\lesssim\|f\|_{\dot{C}^{0,\omega}(B(0,2R))}\omega(\|\tau(x)-\tau(z)\|)\lesssim\omega(\|\tau(x)-\tau(z)\|).\end{split}$ (24)
Let $\theta_{x,z}(M)$ be defined by the identity
$\|\tau(x)-\tau(z)\|=:\theta_{x,z}(M)\|x-z\|,\qquad\theta_{x,z}(M)\lesssim R/\sqrt{M},$
where the bound follows by (18) and (20). By the condition $\lim_{t\to 0}\omega(t)=0$, we choose $M=M(\varepsilon,r,R)$ sufficiently large so that $\omega(RM^{-1/4})\leq\varepsilon\omega(r)$. Since we are assuming $r\leq\|x-z\|\leq M^{1/4},$ we obtain
@@
Chaining the estimates (24) and (25), we have shown (22) for $x,z$ as in the present case.
Finally we consider the case $\tau(x),\tau(z)\in B(0,R),$ and $x\in B(0,2M)\setminus B(0,M)$ and $z\in X\setminus B(0,10M).$ We pick an auxiliary point $y$ such that $|y|=2M.$ Using $\tau(z)=\tau(y)=0,$ the bound (25) (valid by $x,y\in B(0,10M)$) and that $\omega$ is non-decreasing we obtain
$\begin{split}|g(x)-g(z)|=|g(x)-g(y)|\leq\|f\|_{\dot{C}^{0,\omega}(X)}\omega(\|\tau(x)-\tau(y)\|)\lesssim\varepsilon\omega(\|x-y\|)\leq\varepsilon\omega(\|x-z\|),\end{split}$
which is a bound of the correct form. The case $\tau(x),\tau(z)\in B(0,R)$ and $z\in B(0,2M)$ and $x\in X\setminus B(0,10M)$ is symmetrical to the last one.
We now fix the parameters $M\gg R\gg r$ so that both (21) and (22) hold. Then, we show the right claim on the line (15). If $x,z\in B(0,M)$ there is nothing to prove, since $\tau(x)=x$ and $\tau(z)=z.$ Finally, we check that
$\sup_{x\in B(0,M)^{c}}\sup_{z\in\mathbb{R}^{n}}\frac{\left|(f-g)(x)-(f-g)(z)\right|}{\omega(\|x-z\|)}$
$\qquad\leq\sup_{x\in B(0,M)^{c}}\sup_{z\in X}\frac{\left|f(x)-f(z)\right|}{\omega(\|x-z\|)}+\sup_{x\in B(0,M)^{c}}\sup_{z\in X}\frac{\left|g(x)-g(z)\right|}{\omega(\|x-z\|)}\lesssim\varepsilon,$
and hence have now shown both claims on the line (15). ∎

### v4

@@ 1.5. Applications to the compactness of the bi-commutator
The product BMO upper bound for the $L^{p}$-to-$L^{p}$ boundedness of the bi-commutator was obtained with the Hilbert transforms in Ferguson, Sadosky [FerSad] and in full generality later by Dalenc, Ou [DaO16]. The proof of the corresponding lower bound in terms of the product $\mathrm{BMO}$ for specific singular integrals (Hilbert/Riesz transforms) has been reported to have a gap, see e.g. [Ferguson2002, HTV2021ce, AHLMO2021], and a fix has not yet been published. The existence of this gap adds considerable interest to all natural questions involving the bi-commutator.
@@
Instead of looking at $L^{p}$-to-$L^{p}$ bounds, the product stucture of the bi-commutator makes it natural to consider mixed $L^{p_{1}}(L^{p_{2}})$-to-$L^{q_{1}}(L^{q_{2}})$ bounds, where we allow $q_{i}\not=p_{i}.$ This was done in [AHLMO2021]. The exponents directly connecting with our work are those $p_{i},q_{i}\in(1,\infty)$ for which $\beta_{i}:=n_{i}(1/p_{i}-1/q_{i})\geq 0$ and moreover the inequality $\beta_{i}>0$ is strict for some $i=1,2,$ which we start to assume here and below. The function space to characterize the $L^{p_{1}}(L^{p_{2}})$-to-$L^{q_{1}}(L^{q_{2}})$ bounds turned out to be
@@
$\|b\|_{\operatorname{BMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n})}\\ =\sup_{R=I_{1}\times I_{2}}\ell(I_{1})^{-\beta_{1}}\fint_{I_{1}}\ell(I_{2})^{-\beta_{2}}\fint_{I_{2}}|b(x_{1},x_{2})-\langle b(x_{1},\cdot)\rangle_{I_{2}}-\langle b(\cdot,x_{2})\rangle_{I_{1}}+\langle b\rangle_{I_{1}\times I_{2}}|\,\mathrm{d}x_{2}\,\mathrm{d}x_{1}.$
@@
The space $\operatorname{BMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n})$ is in fact a $\mathrm{BMO}$-valued or a $\dot{C}^{0,\alpha}$-valued Hölder space, see [AHLMO2021, Section 3]. Then, recently in [MarOik24], the corresponding $L^{p_{1}}(L^{p_{2}})$-to-$L^{q_{1}}(L^{q_{2}})$ compactness was characterized through the VMO analogue $\operatorname{VMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n}).$ Interestingly, such mixed norm compactness results can be used to prove the sufficiency of the natural product VMO condition for compactness in the non-mixed $L^{p}\to L^{p}$ case. That proof rests on an adaptation of an extrapolation scheme for commutators from [Oik25] and the following Theorem 1.11, which is a corollary of the approximations developed in this article.
@@
Let $\beta_{i}\in[0,1)$ and $\beta_{i}>0$ for some $i=1,2.$ Then,
@@
Theorem 1.11 is a bi-parameter analogue of Theorem 1.10. The proof of Theorem 1.11 is split between the current article and [MarOik24]. We record a quantitative version of Theorem 1.6 in the body-text after the proof of Theorem 1.6 as Corollary 2.5, and then, assuming Corollary 2.5, Theorem 1.11 is proved in [MarOik24, Section 3.].
@@ Acknowledgements
We thank the anonymous referee for thoroughly reading the manuscript and for their comments, which led to improvements in both clarity and presentation. C. M. was supported by the grant no. 334466 of the Research Council of Norway, “Fourier Methods and Multiplicative Analysis”. T.O. was supported by the Finnish Academy of Science and Letters, and by the MICINN (Spain) grant no. PID2020-114167GB-I00.
@@ 2. Bounded support approximation
In this section we prove Theorem 1.2 and then show how to obtain from this Theorem 1.6 and then Corollary 2.5. After these, at the end of this section, we prove Theorem 1.9. Recall that we simplify the notation by denoting $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|.$ We denote by $B(x,r):=\{z\in X\,:\,\|x-z\|\leq r\}$ the closed ball with center $x\in X$ and radius $r>0.$ We begin with a simple remark.
@@
First, note that if $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ has bounded support, then, by property (3) of $\omega,$ $f$ belongs to $\dot{\operatorname{VC}}^{0,\omega}(X,Y).$ Since $\dot{\operatorname{VC}}^{0,\omega}(X,Y)$ is closed under limits with respect to the seminorm $\dot{C}^{0,\omega},$ as per Remark 2.1, the inclusion $\subset$" holds true.
The reverse inclusion is much more complicated. To construct the desired approximation, we define, for each $M>0,$ the following key function
$\tau_{M}:X\to B(0,M),\qquad\tau_{M}(x)=\begin{cases}x,\qquad&\|x\|<M,\\ \left(\frac{2M-\|x\|}{M}\right)^{2}x,\qquad M\leq&\|x\|<2M,\\ 0,\qquad&\|x\|\geq 2M.\end{cases}$ (12)
Our goal is to show that, given $f\in\dot{\operatorname{VC}}^{0,\omega}$ and $\varepsilon>0,$ we can find $M>0$ large enough so that
$f\circ\tau_{M}\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}},\qquad\|f-f\circ\tau_{M}\|_{\dot{C}^{0,\omega}}\lesssim\varepsilon.$ (13)
Notice that $f\circ\tau_{M}-f(0)$ is zero outside $B(0,2M),$ and hence $f\circ\tau_{M}$ has bounded support. (Recall that $\dot{C}^{0,\omega}$ is defined modulo constant equivalence classes.) The choice of $M$ will be made in terms of a small parameter $r$ and a large parameter $R,$ both depending on $\varepsilon$ and $f.$ But before doing that, we need to check some properties concerning the function $\tau_{M}$, for any $M>0.$
First, we verify that
$\|\tau_{M}(x)-\tau_{M}(z)\|\leq 5\|x-z\|\quad\text{for all}\quad x,z\in X.$ (14)
Indeed, for points $x,z\in B(0,2M)\setminus B(0,M),$ the definition of $\tau$ and the triangle inequality give
$\begin{split}\|\tau_{M}(x)-\tau_{M}(z)\|&\leq\Big|\big(2-\frac{\|x\|}{M}\big)^{2}-\big(2-\frac{\|z\|}{M}\big)^{2}\Big|\|z\|+\big(2-\frac{\|x\|}{M}\big)^{2}\|x-z\|\\ &\leq\big(2-\frac{\|x\|}{M}+2-\frac{\|z\|}{M}\big)\frac{\|x-z\|}{M}\|z\|+\big(2-\frac{\|x\|}{M}\big)^{2}\|x-z\|\\ &\leq\big(4-\frac{2M}{M}\big)\frac{\|x-z\|}{M}(2M)+\big(2-\frac{M}{M}\big)^{2}\|x-z\|=5\|x-z\|,\end{split}$ (15)
where in the second bound we used the basic identity $a^{2}-b^{2}=(a+b)(a-b).$ As $\tau_{M}$ is obviously $1$-Lipschitz in the sets $B(0,M)$ and $B(0,2M)^{c},$ and $\tau_{M}$ is continuous in $X,$ we deduce (14) for all $x,z\in X.$
Then we show the following contraction property [^1] If the power $2$ in the definition (12) of $\tau_{M}$ was replaced with the power $1,$ neither (16) nor a variation of (16) with a function $\theta:(0,\infty)\to(0,\infty)$ satisfying $\lim_{M\to\infty}\theta(M)=0$ instead of $1/\sqrt{M}$ is true.:
$\Big\{M>R\geq 1,\quad x\in B(0,M)^{c},\quad\tau_{M}(x),\tau_{M}(z)\in B(0,R)\Big\}\\ \Longrightarrow\|\tau_{M}(x)-\tau_{M}(z)\|\leq\frac{5R}{\sqrt{M}}\|x-z\|.$ (16)
To show (16), assume $M>R\geq 1,$ and we distinguish two cases. If $\|x-z\|>M/2,$ then we have
$\|\tau_{M}(x)-\tau_{M}(z)\|\leq 2R<2R\frac{\|x-z\|}{M/2}=\frac{4R}{M}\|x-z\|\leq\frac{5R}{\sqrt{M}}\|x-z\|.$
So next assume that $z\in B(x,M/2).$ Since $R<M$ and $\tau_{M}(z)\in B(0,R)$, by the definition of $\tau$ we must have necessarily $\|z\|>M.$ Thus, it remains to show that
$\Big\{M>R\geq 1,\quad x,z\in B(0,M)^{c},\quad\tau_{M}(x),\tau_{M}(z)\in B(0,R)\Big\}\\ \Longrightarrow\|\tau_{M}(x)-\tau_{M}(z)\|\leq\frac{5R}{\sqrt{M}}\|x-z\|,$ (17)
which is symmetric with respect to both variables $x,z.$ Notice that if $x,z\in B(0,2M)^{c},$ then the left-hand side of the claimed estimate is zero. Assume that $x,z\in B(0,2M)\setminus B(0,M)$ and we make the following observation. By $\tau_{M}(x)\in B(0,R)$ and $x\in B(0,M)^{c},$ there holds that
$\Big\|\left(\frac{2M-\|x\|}{M}\right)^{2}x\Big\|\leq R,\quad\text{and so}\quad\Big|\frac{2M-\|x\|}{M}\Big|\leq\sqrt{\frac{R}{\|x\|}}\leq\sqrt{\frac{R}{M}},$
and the same bound is valid for the variable $z.$ Thus, by (15), and also using $M>R\geq 1$, we obtain
$\|\tau_{M}(x)-\tau_{M}(z)\|$ $\leq\Big(2-\frac{\|x\|}{M}+2-\frac{\|z\|}{M}\Big)\frac{\|x-z\|}{M}\|z\|+\Big(2-\frac{\|x\|}{M}\Big)^{2}\|x-z\|$
$\leq\Big(4\sqrt{\frac{R}{M}}+\frac{R}{M}\Big)\|x-z\|\leq 5\sqrt{\frac{R}{M}}\|x-z\|\leq\frac{5R}{\sqrt{M}}\|x-z\|.$
As the last case, suppose that $z\in B(0,2M)^{c}.$ This case follows by the continuity of $\tau_{M}$ and the previous case. Indeed, taking a point $y\in[x,z]$ with $\|y\|=2M,$ by the previous case applied for $x$ and $y,$ and bearing in mind that $\tau_{M}(z)=\tau_{M}(y)=0,$ we deduce
$\|\tau_{M}(x)-\tau_{M}(z)\|=\|\tau_{M}(x)-\tau_{M}(y)\|\leq\frac{5R}{\sqrt{M}}\|x-y\|\leq\frac{5R}{\sqrt{M}}\|x-z\|.$
We have shown (17), and thus the claim (16).
We now begin the verification of (13), so let $f\in\dot{\operatorname{VC}}^{0,\omega}$. Since $\tau_{M}$ is Lipschitz (14), Lemma 2.3 implies the left claim on the line (13), for any $M>0.$ Given $\varepsilon>0,$ it remains to determine the parameter $M$ so that the right claim in (13) holds.
Since $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}$, by Lemma 2.3 (and its proof), we take $r>0$, depending on the data $f,\varepsilon,C_{\operatorname{db}}$ (but not on $\sigma$) so that
$\sup_{\begin{subarray}{c}x\not=y\in X\\ \|x-y\|<r\end{subarray}}\frac{|f\circ\sigma(x)-f\circ\sigma(y)|}{\omega(\|x-y\|)}\leq\varepsilon,\quad\text{for every function }\>\sigma:X\to X\text{ with }\operatorname{Lip}(\sigma)\leq 5.$ (18)
Because $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{far}}$, Lemma 2.2 allows us to take large $R>1$ (depending only on $\varepsilon$ and $f$) so that
@@
Finally, by the assumption (2) on $\omega$, we take $M>0$ large enough so that
$M>R,\quad\omega(R\cdot M^{-1/4})\leq\varepsilon\omega(r),\quad\text{and}\quad\omega(2R)\leq\varepsilon\omega(M^{1/4}).$ (20)
In order to prove that (13) holds with this (20) choice of $M>0,$ we will first show that
@@
Fix a point $x\in B(0,M)^{c}$ and we distinguish cases depending on the sizes of $\tau(x)$ and $\tau(z).$
Case 1. If $\tau_{M}(x)\in B(0,R)^{c},$ then by (19), (14), and $\omega$ being non-decreasing and doubling (4),
$\left|f(\tau_{M}(x))-f(\tau_{M}(z))\right|\leq\Big(\sup_{u\in B(0,R)^{c}}\sup_{v\in X}\frac{\left|f(u)-f(v)\right|}{\omega(\|u-v\|)}\Big)\omega(\|\tau_{M}(x)-\tau_{M}(z)\|)\lesssim_{C_{\operatorname{db}}}\varepsilon\omega(\|x-z\|).$
Case 2. The case $\tau_{M}(z)\in B(0,R)^{c}$ is symmetrical to Case 1.
Case 3. Now consider $\tau_{M}(x),\tau_{M}(z)\in B(0,R)$. First observe that when $\|x-z\|\geq M^{1/4},$ using that $f\in\dot{C}^{0,\omega}$ and (20), we have
$\left|f(\tau_{M}(x))-f(\tau_{M}(z))\right|\leq\|f\|_{\dot{C}^{0,\omega}}\omega(|\tau_{M}(x)-\tau_{M}(z)|)\\ \leq\|f\|_{\dot{C}^{0,\omega}}\omega(2R)\leq\|f\|_{\dot{C}^{0,\omega}}\varepsilon\omega(M^{1/4})\leq\|f\|_{\dot{C}^{0,\omega}}\varepsilon\omega(\|x-z\|).$
Therefore, as the second subcase of the current Case 3, we now assume that $\|x-z\|\leq M^{1/4}$. Moreover, (18) and the fact that $\tau_{M}$ is $5$-Lipschitz (see (14)) imply that, to prove the estimate (21) for $x,z,$ we may assume that $\|x-z\|\geq r$. Thus, our current standing assumptions are
$x\in B(0,M)^{c},\quad\tau_{M}(x),\tau_{M}(z)\in B(0,R),\quad r\leq\|x-z\|\leq M^{1/4}.$ (22)
First we bound
$\left|f(\tau_{M}(x))-f(\tau_{M}(z))\right|\lesssim\|f\|_{\dot{C}^{0,\omega}(B(0,R))}\omega(\|\tau_{M}(x)-\tau_{M}(z)\|)\lesssim_{f}\omega(\|\tau_{M}(x)-\tau_{M}(z)\|).$ (23)
Then, by the property (16) (first two of (22) are in force), the upper bound $\|x-z\|\leq M^{1/4}$ of (22), then the second inequality of (20), and finally the lower bound $r\leq\|x-z\|$ of (22) give us
[…]
where we used also that $\omega$ is non-decreasing. Chaining the estimates (23) and (24), we have shown (21) for $x,z$ as in the present Case 3.
We are finally ready so show the right claim on the line (13). If $x,z\in B(0,M)$ there is nothing to prove, since $\tau(x)=x$ and $\tau(z)=z.$ The supremum over those $x\in B(0,M)^{c},$ $z\in X,$ can be estimated using (19), the fact that $M>R$ (see (20)), and (21) as
$\sup_{x\in B(0,M)^{c}}\sup_{z\in X}\frac{\left|(f-(f\circ\tau_{M}))(x)-(f-(f\circ\tau_{M}))(z)\right|}{\omega(\|x-z\|)}$
$\qquad\leq\sup_{x\in B(0,M)^{c}}\sup_{z\in X}\frac{\left|f(x)-f(z)\right|}{\omega(\|x-z\|)}+\sup_{x\in B(0,M)^{c}}\sup_{z\in X}\frac{\left|(f\circ\tau_{M})(x)-(f\circ\tau_{M})(z)\right|}{\omega(\|x-z\|)}$
$\qquad\lesssim\sup_{x\in B(0,R)^{c}}\sup_{z\in X}\frac{\left|f(x)-f(z)\right|}{\omega(\|x-z\|)}+\varepsilon\lesssim\varepsilon.$
We have now shown both claims on the line (13), thus completing the proof of Theorem 1.2. ∎